OpenStudy (anonymous):

tanA+tanB/1-tanA.tanB=sin(a+b)/cos(a+b)

sam (.sam.):

You might wanna use sum to product formula.

terenzreignz (terenzreignz):

Proving identities?

OpenStudy (anonymous):

hy sam

OpenStudy (anonymous):

can u solve my optional math question

terenzreignz (terenzreignz):

Let's start with the right-hand side \[\Large \frac{\sin(A+B)}{\cos(A+B)}\] Now hopefully, recalling the sum identitity \[\Large = \frac{\sin(A)\cos(B) + \cos(A)\sin(B)}{\cos(A)\cos(B) - \sin(A)\sin(B)}\]

terenzreignz (terenzreignz):

You there, @bijay56 ? Catch me so far?

OpenStudy (anonymous):

hy terenzreignz

terenzreignz (terenzreignz):

You can call me Terence :) More importantly, do you understand the expansion?

OpenStudy (anonymous):

yes i uderstood and after that

terenzreignz (terenzreignz):

After that, a tricky step... we'll multiply 1 to the expression... but not just any "1" but a special variant of 1... \[\Large = \frac{\sin(A)\cos(B) + \cos(A)\sin(B)}{\cos(A)\cos(B) - \sin(A)\sin(B)}\times \frac{\frac{1}{\cos(A)\cos(B)}}{\frac1{\cos(A)\cos(B)}}\]

terenzreignz (terenzreignz):

So, going ahead and distributing... \[\Large = \frac{\frac{\sin(A)\cos(B)}{\cos(A)\cos(B)} + \frac{\cos(A)\sin(B)}{\cos(A)\cos(B)}}{\frac{\cos(A)\cos(B)}{\cos(A)\cos(B)} -\frac{ \sin(A)\sin(B)}{\cos(A)\cos(B)}}\]

terenzreignz (terenzreignz):

The fun part... I'll leave this to you, but cancel out what can be cancelled, and remember that... \[\huge \frac{\sin(x)}{\cos(x)}=\tan(x)\] Enjoy the slashing ;)

OpenStudy (anonymous):

thanks terence

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

i have some work i will come in minute ternce

terenzreignz (terenzreignz):

\[\huge = \frac{\frac{\sin(A)\cos(B)}{\cos(A)\cos(B)} + \frac{\cos(A)\sin(B)}{\cos(A)\cos(B)}}{\frac{\cos(A)\cos(B)}{\cos(A)\cos(B)} -\frac{ \sin(A)\sin(B)}{\cos(A)\cos(B)}}\] this is almost done, by the way, just some slashing and converting, and you'll be able to turn this into the left-side in a jiffy.

OpenStudy (anonymous):

hy

terenzreignz (terenzreignz):

Yes?

OpenStudy (anonymous):

cos(a+b).cos(a-b)=cos^2A-sin^2B

terenzreignz (terenzreignz):

Expand the left-hand side.

OpenStudy (anonymous):

how

terenzreignz (terenzreignz):

This rule... \[\large \cos(\alpha+\beta) = \cos(\alpha)\cos(\beta) - \sin(\alpha)\sin(\beta)\]

OpenStudy (anonymous):

ok and after that

terenzreignz (terenzreignz):

Expand the left-hand side first, then I'll tell you the next step.

OpenStudy (anonymous):

i have done now

terenzreignz (terenzreignz):

What did you get?

OpenStudy (anonymous):

cosA.cosB-sinA.sinB.cosA.cosB+sinA.sinB

terenzreignz (terenzreignz):

Okay. But put the parentheses. (cosAcosB - sinAsinB)(cosAcosB + sinAsinB) Notice that it's of the form... \[\large (x-y)(x+y)\]which is equal to... \[\large x^2 - y ^2\]

OpenStudy (anonymous):

now what to do

OpenStudy (anonymous):

terence now what to do after that

terenzreignz (terenzreignz):

What to do? I told you already... \[\Large [\cos(A)\cos(B) -\sin(A)\sin(B)][\cos(A)\cos(B) + \sin(A)\sin(B)]\] It's just like this... \[\Large (x-y)(x+y) = x^2 - y^2\]

terenzreignz (terenzreignz):

Sorry, got cropped...\[\large [\cos(A)\cos(B) -\sin(A)\sin(B)][\cos(A)\cos(B) + \sin(A)\sin(B)]\]

terenzreignz (terenzreignz):

It's like... \[\large x = \cos(A)\cos(B) \\\large y = \sin(A)\sin(B)\]

OpenStudy (anonymous):

now next step

terenzreignz (terenzreignz):

You know me :P No next step until you show me your progress :D

OpenStudy (anonymous):

next step:cos^2A.cos^2b-sin^2A,sin^2B

OpenStudy (anonymous):

next step:cos^2A.cos^2B-sin^2A.sin^2B

OpenStudy (anonymous):

next step right or wrong

terenzreignz (terenzreignz):

And from here, it's just algebraic manipulation.

terenzreignz (terenzreignz):

Whoops, sorry.... \[\large \cos^2(A)\cos^2(B)- \color{green}{\sin^2(A)}\sin^2(B)\] replace with...\[\large \cos^2(A)\cos^2(B)- \color{red}{[1-\cos^2(A)]}\sin^2(B)\]

terenzreignz (terenzreignz):

If it makes it simpler, it may be tweaked into this... \[\large \cos^2(A)\cos^2(B) \color{blue}{+[\cos^2(A)-1]}\sin^2(B)\]

OpenStudy (anonymous):

hy

OpenStudy (anonymous):

terence can u solve me another opt math problem